OpenStudy (anonymous):

solve and write as interval notation 5/9 < x -1/9 < 6

OpenStudy (lgbasallote):

\[\huge \frac 59 < x - \frac 19 < 6\] firs step is to add 1/9 to both sides...

OpenStudy (anonymous):

so 5/9<x<1/9+6?

OpenStudy (lgbasallote):

when i said both.. i meant all sides

OpenStudy (lgbasallote):

\[\huge \frac 59 + \frac 19 <x < \frac 19 + 6\] now simplify this

OpenStudy (anonymous):

so 1/9+5/9<x<1/9+6?

OpenStudy (lgbasallote):

yes. now simplify it

OpenStudy (lgbasallote):

combine the like terms

OpenStudy (anonymous):

Add 1/9 to the each part, then reduce the parts that you can. After doing so, you get the following: \[\frac{ 2 }{ 3 }<x<\frac{ 55 }{ 9 }\]

OpenStudy (anonymous):

2/3<x<55/9 is the answer in interval notation

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

That is what you get when you add 1/9 to each part and then reduce each part to the simplest form.

OpenStudy (unklerhaukus):

\[(2/3,55/9)\]